OpenStudy (anonymous):

2CH6 + 7O2 --> 4CO2 + 6H2O When I already have the balanced equation, how do I find how many grams of CO2 are produced when 2.00 moles of C2H6 react with sufficient O2? I'm having trouble understanding finding how many grams needed for a reaction with grams/moles of another substance.

OpenStudy (mos1635):

2C2H6 + 7O2 --> 4CO2 + 6H2O 2moles of C2H6 produces 4 moles of CO2 n=m/Mr......

OpenStudy (anonymous):

It asks how many grams of CO2...

thomaster (thomaster):

You can calculate the amount of grams if you know the amount of moles. To do this you need the molecular weight of the compound. You will need a periodic table to calculate the molecular weight. First check what the compound is made of. In this case CO2 consists of carbon and oxygen Look on a periodic table and you'll see: 1 mole of carbon weights 12.0107 grams 1 mole of oxygen weights 15.9994 grams You have 1 carbon and 2 oxygen so 12.0107 + (2*15.9994) = 44.0095 gram/mole So 1 mole CO2 weights 44.0095 grams You have 4 mole CO2 so just 4* 44.0095 = 176.0380 gram CO2

thomaster (thomaster):

oh and a nice periodic table : http://ptable.com